OpenStudy (anonymous):

factors of 6 that adds up to 17???

OpenStudy (sriram):

x^2-17x+6=0

OpenStudy (sriram):

roots are 0.4 and 16.2 approx

OpenStudy (sriram):

these roots when added gives 17 and multiplied 6

OpenStudy (anonymous):

how do you find that @sririam? Quadratic equation?

OpenStudy (sriram):

in the quadreatic ax^2+bx+c sum of roots =-b/a and their product=c/a hence we want 2 numbers whose product=6 and sum=17 i took a=1 then -b=17 & c=6

OpenStudy (anonymous):

storing away that piece of knowledge :-)... thanks!

OpenStudy (anonymous):

No wait- I mean how did you get the 0.4 and 16.2